# Low-Rank Tensor Decomposition over Finite Fields

### v3

@@ 2.3 Runtime
Given a matrix $M\in\mathbb{F}^{m\times n}$ and integer $r^{\prime}$, we can do the following in $O(mn\mathrm{min}\left(r^{\prime},\mathrm{rank}\left(M\right)\right))$ time:
@@
Finally, each execution of step (b) involves $O(m)$ many row operations in $F$ and column operations in $C$, each of which takes $O(n)$ time, so the total runtime is $O(mn\mathrm{min}\left(r^{\prime},\mathrm{rank}\left(M\right)\right))$. 777 As a sidenote, we can slightly modify this algorithm to solve Lemma 1 in $O(mn\mathrm{rank}\left(M\right))$ time: remove the early termination (ix.); maintain another matrix $T$ that is initialized to $I_{m}$, and perform the same operations on it as the ones we do to $F$ throughout the algorithm;
In our algorithm, we first extract a basis of tensor slices $T_{i,:,:}$ by row-reducing the $n\times n^{2}$ matrix $\widetilde{T}:=\begin{bmatrix}\vdots\\ \hline\cr\overrightarrow{T_{i,:,:}}\\ \hline\cr\vdots\end{bmatrix}_{i}$. We also check whether $\mathrm{rank}\left(\widetilde{T}\right)\leq R$ and immediately exit if this condition is not satisfied. Using Lemma 5, we can complete this entire step in $O(Rn^{3})$ time.
@@
To solve the system over $M_{r}$, we first separate it into independent parts, of which there are $\leq R$ many. Then we solve each part by enumerating full-rank factorizations of a matrix with rank $\leq g(R)$, for some function $g$, and compute $O(R)$ many additional rank factorizations to extract $B_{r,:}$ and $C_{r,:}$ if a solution is found.
To figure out $g$:
  • If $R\leq 2$: all coefficient matrices $E$ are separable into single rows, so we do not enumerate full-rank factorizations. For this, we define $g(R):=0$.
@@
  • If $R=3$: the most difficult case occurs when $E=\begin{bmatrix}1&&x\\ &1&y\end{bmatrix}$ and at least one of $D_{0}$ or $D_{1}$ has rank 2.
  • If $R=4$: the most difficult case occurs when $E=\begin{bmatrix}1&&w&x\\ &1&y&z\end{bmatrix}$ and at least one of $\Phi:=\left[\begin{array}[]{c|c}D_{0}&D_{1}\end{array}\right]$ or $\Theta:=\left[\begin{array}[]{c}D_{0}\\ \hline\cr D_{1}\end{array}\right]$ has rank 4.
    • We also have the subcase where $\mathrm{rank}\left(D_{0}\right)=\mathrm{rank}\left(D_{1}\right)=2\ \land\ \mathrm{rank}\left(\Phi\right)=\mathrm{rank}\left(\Theta\right)=3\ \land\ \mathrm{rank}\left(U\right)=3$, where we enumerate every $3\times 4$ matrix $F$ and solve for row-vectors $v_{i}$ in a linear system. Since this system contains $O(n)$ total scalar variables, it can be calculated in $O(n^{3})$ time, so this entire subcase can be solved in $O(|\mathbb{F}|^{12}n^{3})$ time, which is asymptotically less costly than the case where at least one of $\Phi$ or $\Theta$ has rank 4.
Thus, we have $g(1)=0,\ g(2)=0,\ g(3)=2,\ g(4)=4$.
Since all rank factorizations are done for $O(n)\times O(n)$ matrices, computing each one takes $O(n^{3})$ time, so we can upper bound the total time complexity of our algorithm at $O\left(Rn^{3}+|\mathbb{F}|^{R^{2}}R\left(|\mathbb{F}|^{g(R)^{2}}+R\right)n^{3}\right)=O\left(\left(R+|\mathbb{F}|^{R^{2}}R\left(|\mathbb{F}|^{g(R)^{2}}+R\right)\right)n^{3}\right)$, which is sufficient to prove Theorem 1.
2.3.1 Tighter analysis
By analyzing the rank factorizations of our algorithm more carefully, we can obtain a tighter bound.
Consider enumerating full-rank factorizations for a $O(n)\times O(n)$ matrix $M$ with rank $r$. We first compute an initial factorization $M=CF$, where $C$ has $r$ columns and $F$ has $r$ rows, which can be done in $O(rn^{2})$ time. Then, for each $X\in\mathrm{GL}\left(r,\ \mathbb{F}\right)$, we compute $(CX,X^{-1}F)$, which takes $O(nr^{2})$ time (we can precompute each $X^{-1}$ ahead of time). However, we may also have to set each appropriate variable matrix $M_{r}$ to a summand $(CX)_{:,r^{\prime}}(X^{-1}F)_{r^{\prime},:}$ of the rank factorization, e.g., in the rows-sharing-one-column case, in order to continue with the rest of the algorithm, so constructing these $M_{r}$ will take additional $O(rn^{2})$ time. Thus, the total time needed to enumerate all full-rank factorizations and set the appropriate $M_{r}$ is $O(|\mathbb{F}|^{r^{2}}(rn^{2}+r^{2}n))$.
Lastly, we must address the one subcase under $R=4$ that we mentioned earlier: where $\mathrm{rank}\left(D_{0}\right)=\mathrm{rank}\left(D_{1}\right)=2\ \land\ \mathrm{rank}\left(\Phi\right)=\mathrm{rank}\left(\Theta\right)=3\ \land\ \mathrm{rank}\left(U\right)=3$, and we must solve $F\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]=W_{*}$, where $F\in\mathbb{F}^{3\times 4}$, $v_{i}\in\mathbb{F}^{1\times n}$, and $W_{*}\in\mathbb{F}^{3\times 2n}$ forms the right side of a rank-3 factorization of $\Phi\in\mathbb{F}^{n\times 2n}$ (i.e. $\Phi=U_{*}W_{*}$ for some $U_{*}\in\mathbb{F}^{n\times 3}$). We can actually separate this system into $n$ many disjoint linear systems $F\begin{bmatrix}v_{0,i}&0\\ 0&v_{1,i}\\ v_{2,i}&v_{2,i}\\ v_{3,i}&zv_{3,i}\end{bmatrix}=\left[\begin{array}[]{c|c}W_{:,i}&W_{:,n+i}\end{array}\right]$ for each $0\leq i<n$, where each system only consists of $8$ variables instead of $O(n)$. After accounting for the enumeration of $F$ and the factorization $\Phi=U_{*}W_{*}$, we can actually solve this entire subcase in $O(3n^{2}+|\mathbb{F}|^{12}8^{3}n)$ time; this is less than the other subcase under $R=4$ where $\mathrm{rank}\left(\Phi\right)=4\ \lor\ \mathrm{rank}\left(\Theta\right)=4$, which requires $O(|\mathbb{F}|^{16}(4n^{2}+4^{2}n))$ time.
Applying these findings to our algorithm, we enumerate rank factorizations of a $O(n)\times O(n)$ matrix of rank $g(R)$; and if a solution is found, we rank-factor the $D_{i}$ matrices, each of which must have rank $\leq R$ by the rank precondition (or else no solution exists and we terminate), which takes $O(R)\cdot O(Rn^{2})=O(R^{2}n^{2})$ additional time.
Thus, the time complexity of our algorithm is $O\left(Rn^{3}+|\mathbb{F}|^{R^{2}}R\left(|\mathbb{F}|^{g(R)^{2}}\left(g(R)n^{2}+g(R)^{2}n\right)+R^{2}n^{2}\right)\right)$. Notice that the $n^{3}$ now has a small coefficient compared to the other terms. It may be possible to obtain slightly tighter bounds through more careful analysis.

### v4

@@ 2.3 Runtime
Since $R\leq 4$, we will omit $R$ terms in some places to simplify analysis.
@@
Given a matrix $M\in\mathbb{F}^{m\times n}$ and integer $r^{\prime}$, we can do the following in $O(m^{2}+m(m+n)\mathrm{min}\left(r^{\prime},\mathrm{rank}\left(M\right)\right))$ time:
@@
Finally, each execution of step (b) involves $O(m)$ many row operations in $F$ and column operations in $C$, each of which takes $O(m+n)$ time. Since initializing $T$ takes $O(m^{2})$ time, the total runtime is $O(m^{2}+m(m+n)\mathrm{min}\left(r^{\prime},\mathrm{rank}\left(M\right)\right))$. ∎
Throughout our algorithm for tensor decomposition, we maintain the invariants that $T$ is invertible and $TM=F$. At the end of the algorithm, $F=\mathrm{rref}\left(M\right)$.
In our algorithm, we first extract a basis of tensor slices $T_{i,:,:}$ by row-reducing the $n\times n^{2}$ matrix $\widetilde{T}:=\left[\begin{array}[]{c}\vdots\\ \hline\cr\overrightarrow{T_{i,:,:}}\\ \hline\cr\vdots\end{array}\right]_{i}$. We also check whether $\mathrm{rank}\left(\widetilde{T}\right)\leq R$ and immediately exit if this condition is not satisfied. Using Lemma 6, we can complete this entire step in $O(Rn^{3})$ time, instead of the more naïve $O(n^{4})$ time.
@@
To solve the system over $M_{r}$, we first separate it into independent parts, of which there are $\leq R$ many. To solve each part, we consider the most difficult subcases for each $R$:
  • If $R\leq 2$: all coefficient matrices $E$ are separable into single rows, and we must have $\mathrm{rank}\left(D_{i}\right)\leq R$ in order for a solution to exist, so we can find a solution (or prove it does not exist) in $O(n^{2})$ time.
  • If $R=3$: the most difficult case occurs when $E=\begin{bmatrix}1&&x\\ &1&y\end{bmatrix}$ and at least one of $D_{0}$ or $D_{1}$ has rank 2, where we have to enumerate $O(|\mathbb{F}|^{4})$ many full-rank factorizations of such a $D_{i}$. For each factorization of $D_{0}$ (or $D_{1}$), we check if $D_{0}-xM_{2}$ (or $D_{1}-yM_{2}$) has rank $\leq 1$. Overall, solving this case takes $O(|\mathbb{F}|^{4}n^{2})$ time.
@@
  • If $R=4$: the most difficult case occurs when $E=\begin{bmatrix}1&&w&x\\ &1&y&z\end{bmatrix}$ and both $D_{0}$ and $D_{1}$ have rank 2, where we have to enumerate $O(|\mathbb{F}|^{12})$ many pairs of matrices in 4 subcases each.
    Generating each pair of matrices takes $O(n)$ time, and for each pair we must solve for $X$ in some linear system $AX=B$ (or its transpose), where $A$ has shape $O(n)\times O(R)$ and $X$ has shape $O(R)\times O(n)$.
    We can find $X$ with row reduction: find some invertible $T$ s.t. $TA=\mathrm{rref}\left(A\right)$ and extract a solution $X$ from the row-reduced augmented matrix $T\left[\begin{array}[]{c|c}A&B\end{array}\right]=\left[\begin{array}[]{c|c}\mathrm{rref}\left(A\right)&TB\end{array}\right]$, e.g., by checking for all-zeros rows in $\mathrm{rref}\left(A\right)$ and setting free variables of $X$ to 0.
    We can obtain such $T$ in $O(n^{2}R)$ time by modifying the algorithm in Lemma 6: remove early termination, maintain the inverse of the $C$ matrix by performing row operations on a separate matrix, and continue the row reduction so that the $F$ matrix is transformed into rref.
    Overall, solving this subcase takes $O(|\mathbb{F}|^{12}n^{2})$ time.
Thus, we spend $O(|\mathbb{F}|^{g(R)}n^{2})$ time solving each system over $M_{r}$, where $g(1)=0,\ g(2)=0,\ g(3)=4,\ g(4)=12$.
We can then upper bound the total runtime of our algorithm at $O\left(Rn^{3}+|\mathbb{F}|^{R^{2}}R\cdot|\mathbb{F}|^{g(R)}n^{2}\right)$, which for fixed $R$ is equivalent to $O\left(n^{3}+|\mathbb{F}|^{R^{2}+g(R)}n^{2}\right)$. Notice that the large constant is multiplied with $n^{2}$, not $n^{3}$. It may be possible to obtain tighter bounds through more careful analysis.